OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6). 7 −14 0 14

OpenStudy (radar):

\[slope (m)=\Delta y/\Delta x\]

OpenStudy (radar):

\[\Delta y=(8-(-6))=14\]\[\Delta x=(-4-(-3)=-1\] 14/-1 = -14